# The Illusion of Diminishing Returns: Measuring Long Horizon Execution in LLMs

### v2

@@ Acknowledgements
Appendix A Investigating Potential Fixes for Self-conditioning
A.1 Turn-wise Self-Verification Prompting
@@
These findings suggest that prompting self-correction may not be a viable solution. It is computationally expensive, incurring a context-length penalty, and is itself a complex, error-prone execution task that models may not be able to perform reliably.
@@ A.2 Context Engineering
As shown in Figure˜9 (a), performance improves significantly as the context window size is reduced, allowing models to sustain execution for longer horizons. While a fixed sliding window is only applicable to tasks without long-range dependencies, this result validates a more general principle: active context management designed to minimize the accumulation of errors in the context is a promising direction for improving long-horizon reliability in LLM agents.
@@
Appendix B Can Parallel Test-time Compute Scaling Match Thinking?
@@
We also experiment to validate if parallel scaling in test-time compute can achieve the same improvements as thinking. We verify this by testing if parallel majority voting can replicate the gains from either model scale or sequential computation (thinking). To create a fair comparison, we sample multiple outputs from a non-thinking Gemma3 model at each turn, with the number of samples set to match the average token count of its CoT counterpart. The final answer is determined by a majority vote over these parallel generations. From the results in Figure˜10 and 9 (b), we see that while majority voting yields a marginal performance improvement over the base model, it is insufficient to match the reliability of a larger, non-thinking model, let alone the substantial gains from using CoT reasoning. This suggests that for long-horizon execution, sequential computation provides an advantage that parallel test time scaling cannot match. This contrasts with findings in other domains, such as math or common-sense reasoning, where parallel sampling with self-consistency has been shown to be highly competitive (Snell et al., 2024).
Appendix C Number of Turns vs Turn Complexity
In our experiments, we show that we can increase the length of the task needed to be performed by either (1) increasing the number of turns or (2) increasing the turn complexity, i.e, providing more inputs in the same turn. To investigate the relationship between these two axes, we perform an experiment where a model has to perform a fixed number of operations while varying the turn complexity. With a higher turn complexity, the model requires fewer turns to reach the fixed number of operations. Results in Figure˜11 (a) indicate there is no strict turn complexity that is consistently the best across model families. Rather, we found that different models behave quite differently for the same turn complexities. Qwen3 32B seems to show poorer performance at lower turn complexities, indicating that it is unable to perform well over a large number of turns, even if the turns are simple themselves. Gemma3 12B shows a different trend. It reaches accuracy peaks at either extreme of the turn complexity spectrum, failing badly at mid-level turn complexities. This indicates it suffers when the turn complexity and the number of turns are both sufficiently high.
Another axis of evaluating the number of turns vs turn complexity trade-off is the test-time compute used. From a cost perspective, increasing the number of turns increases the overall cost of inference. We can lower the number of turns by increasing the turn complexity, but that would result in an increase in the per-turn inference cost, as a result of the added complexity. For the same experiment above, we track the number of output tokens used for computation (including thinking tokens) per sample, and again find diverging results for each family in Figure˜11 (b).
@@ Appendix C Number of Turns vs Turn Complexity
Figure 12: Scaling trends hold even after enabling Sequential Test Time compute. We compare model performance with thinking disabled (a) against thinking enabled (b, c) at varying turn complexities. (a) Without thinking, all models fail to execute even two steps ($K=2$) in a single turn. (b) In contrast, enabling thinking prevents this performance collapse, with all models successfully handling $K=2$. (c) When the turn complexity is further increased to $K=10$, performance degrades, but a clear scaling trend emerges. (d) This trend is explicitly shown, illustrating that for complex turns, the horizon length increases consistently with model size, reinforcing the benefits of scaling model size even when thinking is enabled.
@@
Appendix D Deconstructing Errors in Retrieve-then-compose Steps
@@
From Figure˜14, we can observe that models achieve near-perfect performance on the stateless retrieval and addition task, indicating that neither simple dictionary lookup nor addition is a significant source of error. In contrast, the prefix sum task, while significantly better than our task, still exhibits a slow degradation over time.
@@
Appendix E Experimental Setup
E.1 Task Details
@@ E.2 Prompting
E.3 Prompting for Thinking Models
@@ E.4 Model Specifications
We use a temperature of $0.6$ and a top-p value of $0.95$ for all Gemma models. For Qwen, we use a temperature of $0.6$ and a top-p value of $0.95$ for thinking mode and a temperature of $0.7$ and a top-p value of $0.8$ for non-thinking as recommended in their documentation. [^1] https://huggingface.co/Qwen/Qwen3-32B We find in Figure˜13 that temperature does not affect the observed trends by much.
E.5 Compute Details
@@
Appendix F Format Following Failures
@@
Our results for format following failures are presented in Figure˜15 for the experiments presented in Section˜3. We observe that smaller models are more susceptible to format failures, with the Qwen3 family in particular being worse at following format instructions. Overall, the fraction for format following errors is low (around 0.1), with the Qwen3-8B being an exception. We find that the error here actually comes from the model trying to cheat, and do the entire summation inside the <answer> tags (For example, <answer>39 + 51 = 90</answer>). This is explicitly forbidden as we do not allow chain-of-thought or thinking in this experiment, and thus we count this as an error. Gemma3 4B fails at later turns due to context length limitations; however, that does not affect any results, as its accuracies drop much earlier.
@@
For the experiments presented in Figure˜15, we find the Qwen3 family to be prone to format following errors in the case where we have thinking disabled for $K=2$. We again find this to be the consequence of models trying to cheat and use extra tokens for computation inside the answer tags. This is fixed by enabling thinking. Following this, the errors in format become negligible. At $K=10$, we see Gemma3 12B sharply rise to a format failure fraction of 1.0, again due to a full context window.
Appendix G Chain-of-Thought Self-Conditioning
@@
The alternative is to programmatically generate flawed CoT traces. We implemented and experimented with this, and as seen in Figure˜16, CoT does not mitigate the self-conditioning effect. However, this setup also introduces its own complexities. For our simple task, there are multiple distinct points of failure within a single trace: an error in the retrieval step (looking up an incorrect value) or an error in the composition step (an arithmetic mistake). A controlled experiment would need to systematically manage the type, frequency, and location of these injected errors, making the setup intractable. Even establishing a “perfectly correct” (Induced Error Rate = 0.00) baseline history is problematic. A model might have a CoT trace with flawed reasoning (e.g., a minor calculation error that cancels out), which we then replace with the correct final answer. Such a history is also unfaithful.
@@
Appendix H Proof and Analysis of Proposition 1
@@
Assuming a constant per-step accuracy $p$ and no self-correction, the horizon-length $H$ at which a model can achieve a success rate $s$ is given by:
@@
H.1 Implications for Horizon Length ($H_{0.5}$)

### v3

@@ Acknowledgements
Appendix A Self-Conditioning in Realistic Agentic Tasks
Zhu et al. (2025a) recently released [^1] Link to AgentErrorBench trajectories annotated failed trajectories from more realistic, popular agent benchmarks like GAIA (Mialon et al., 2024), ALFWorld (Shridhar et al., 2021), and WebShop (Yao et al., 2022). We manually analyzed each annotated sample and identified potential instances of self-conditioning. We attach an example of an annotated entry below.
"trajectory_id": "GPT-4o_001_chat_b000_t00_e00-d3248f80", "LLM": "GPT-4o", "task_type": "webshop", "critical_failure_step": 29, "critical_failure_module": "plan", "step_annotations": [ { "step": 29, "plan": { "failure_type": "inefficient_plan", "reasoning": "This plan loops back to the very beginning, moving in the wrong direction, and results in an inefficient search query that is almost identical to the initial ones." } } ]
We search for entries where the failure reason mentions the agent repeating the previous mistakes repeatedly and eventually degenerating into failure trajectories. Some of the error categories that fit our search were,
  • inefficient_plan: Plan is overly long or illogical because of prior errors.
  • progress_misjudge: Incorrectly evaluates progress based on its own prior outputs.
  • causal_misattribution: Correctly notes failure but blames the wrong cause due to its prior outputs, misguiding subsequent plans.
Through this process, we estimate that roughly $20\%$ of GAIA, $48\%$ of ALFWorld, $33\%$ of WebShop failures are similar to self-conditioning. Here are some such trajectory annotations as examples:
WebShop: “…This set the agent on an endless loop of query reformulation and paging through irrelevant results, making success impossible…”
ALFWorld: “…This set a precedent for all subsequent memory outputs to ignore key details about which objects had been visited what actions were performed at each and what the actual observations were…”
GAIA: “Make an low efficient plan that not success, repeat the similar action that not success”
A caveat in the analysis we present above is that the correctness of steps is subjective to determine on these tasks for a quantitative study of self-conditioning. This is exactly where having a controlled study like ours, even if with a synthetic task, is extremely valuable. It allows us to isolate execution failures from planning and knowledge gaps, which is difficult in more complex agentic tasks.
Appendix B Sequentiality of the Single-Turn Task
A potential limitation of the single-turn setting in Section 3.3 is that retrieving and adding a list of numbers is, in theory, parallelizable inside a transformer (Merrill and Sabharwal, 2023). This is because adding a list of numbers lies in the Threshold Circuit 0 (TC0) (Hajnal et al., 1993) class of problems that can be solved with constant depth computation. While currently our results show that frontier language models fail to perform more than a few such additions without chain of thought, since this is in theory learnable, our single-turn study is not future-proof. Note that this limitation does not apply to our multi-turn experiments, as the keys to be retrieved and their value added is provided turn-by-turn, and thus the model has to perform these steps sequentially. To mitigate this, we now perform a study with a sequence of $2\times 2$ (non-stochastic) matrix multiplications, which in the general case lies in TC1, generally believed to not be equivalent to TC0 (Mereghetti and Palano, 2000).
Experimental Setup.
To keep this task similar to the key-value addition task introduced in Section˜3, we create an analogous key-value matrix multiplication task, where each key corresponds to a randomly generated $2\times 2$ matrix. Since multiplication causes values to blow up to very large numbers quickly, we ensure each element in the matrix is in the range of [9,9][-9,9]. The LLM is then prompted to complete a chain multiplication of $K$ matrices, where $K$ corresponds to the per-turn complexity (Section 3.3). The full prompt is presented below. Analogous to the key-value addition task, we again have a series of retrieve-then-compose steps:
  1. 1.
    Retrieval: Look up the matrix $\mathcal{M}[k]$ for each key $k\in P_{t}$
  2. 2.
    Composition: Compute the product of these matrices and multiply by the previous state, $S_{t}=S_{t-1}\cdot(\prod_{i=1}^{K}\mathcal{M}[k_{t,i}])$
Starting Prompt
You are an AI assistant. I will provide you with a dictionary of keys mapped to 2x2 matrices and give you keys in groups of 2. Maintain the running 2x2 product starting from the identity matrix.
For each group, respond with the current product in row-major order inside <answer></answer> tags using four comma-separated integers.
IMPORTANT: Do not output any other text inside <answer> tags. Do NOT output anything else at all.
Illustrative examples using a reference dictionary:
alpha: [[1, 0], [0, 1]]
beta: [[2, 1], [0, 1]]
gamma: [[1, 0], [1, 1]]
delta: [[0, 1], [1, 0]]
Example 1:
Input Keys: [’alpha’, ’beta’, ’gamma’]
Illustrating with product after each key:
Result: <answer>1,0,0,1 | 2,1,0,1 | 3,1,1,1</answer>
Example 2:
Input Keys: [’delta’, ’beta’, ’gamma’, ’alpha’]
Illustrating with product after each block of 2 keys:
Result: <answer>0,1,2,1 | 1,1,3,1</answer>
Example 3:
Input Keys: [’beta’, ’gamma’, ’beta’]
Illustrating with product after each block of 3 keys:
Result: <answer>6,4,2,2</answer>
Now, here is the actual task:
Dictionary to maintain:
drive: [[-3, 9], [1, -3]]
alone: [[6, 3], [5, -5]]
adult: [[-1, -5], [-2, 8]]...
Ready to start!
IMPORTANT: Only output the four comma-separated integers representing the matrix product inside <answer> tags. Nothing else should be present in the output, ONLY the answer.
Results.
Our findings for the iterated matrix multiplication task, presented in Figure˜8, are similar to those presented in Figure˜7. For each model, we evaluate $50$ independent rollouts. Without sequential test-time compute, models are unable to multiply even 2 matrices. This is expected, as each step of this task requires a constant number of sub-tasks, including multiple retrievals, multiplications, and additions. With sequential test-time compute, we see a substantial increase in each model’s ability to perform the task.
Figure 8: Benchmarking iterative matrix multiplication within a single turn. Without CoT or thinking, even the biggest models fail to multiply just two matrices (a). Sequential test time compute (thinking tokens) significantly improves this (b).
Appendix C Investigating Potential Fixes for Self-conditioning
C.1 Turn-wise Self-Verification Prompting
@@
These findings suggest that prompting-based self-correction may not be enough. It is computationally expensive, incurring a context-length penalty, and is itself a complex, error-prone execution task that models may not be able to perform reliably.
@@ C.2 Context Engineering
As shown in Figure˜10 (a), performance improves significantly as the context window size is reduced, allowing models to sustain execution for longer horizons. While a fixed sliding window is only applicable to tasks without long-range dependencies, this result validates a more general principle: active context management designed to minimize the accumulation of errors in the context is a promising direction for improving long-horizon reliability in LLM agents.
@@
Appendix D Can Parallel Test-time Compute Scaling Match Thinking?
@@
We also experiment to validate if parallel scaling in test-time compute can achieve the same improvements as thinking. We verify this by testing if parallel majority voting can replicate the gains from either model scale or sequential computation (thinking). To create a fair comparison, we sample multiple outputs from a non-thinking Gemma3 model at each turn, with the number of samples set to match the average token count of its CoT counterpart. The final answer is determined by a majority vote over these parallel generations. From the results in Figure˜11 and 10 (b), we see that while majority voting yields a marginal performance improvement over the base model, it is insufficient to match the reliability of a larger, non-thinking model, let alone the substantial gains from using CoT reasoning. This suggests that for long-horizon execution, sequential computation provides an advantage that parallel test time scaling cannot match. This contrasts with findings in other domains, such as math or common-sense reasoning, where parallel sampling with self-consistency has been shown to be highly competitive (Snell et al., 2024).
Appendix E Number of Turns vs Turn Complexity
In our experiments, we show that we can increase the length of the task needed to be performed by either (1) increasing the number of turns or (2) increasing the turn complexity, i.e, providing more inputs in the same turn. To investigate the relationship between these two axes, we perform an experiment where a model has to perform a fixed number of operations while varying the turn complexity. With a higher turn complexity, the model requires fewer turns to reach the fixed number of operations. Results in Figure˜12 (a) indicate there is no strict turn complexity that is consistently the best across model families. Rather, we found that different models behave quite differently for the same turn complexities. Qwen3 32B seems to show poorer performance at lower turn complexities, indicating that it is unable to perform well over a large number of turns, even if the turns are simple themselves. Gemma3 12B shows a different trend. It reaches accuracy peaks at either extreme of the turn complexity spectrum, failing badly at mid-level turn complexities. This indicates it suffers when the turn complexity and the number of turns are both sufficiently high.
Another axis of evaluating the number of turns vs turn complexity trade-off is the test-time compute used. From a cost perspective, increasing the number of turns increases the overall cost of inference. We can lower the number of turns by increasing the turn complexity, but that would result in an increase in the per-turn inference cost, as a result of the added complexity. For the same experiment above, we track the number of output tokens used for computation (including thinking tokens) per sample, and again find diverging results for each family in Figure˜12 (b).
@@ Appendix E Number of Turns vs Turn Complexity
Figure 13: Scaling trends hold even after enabling Sequential Test Time compute. We compare model performance with thinking disabled (a) against thinking enabled for Qwen models and CoT for Gemma models (b, c) at varying turn complexities. (a) Without thinking, all models fail to execute even two steps ($K=2$) in a single turn. (b) In contrast, enabling thinking prevents this performance collapse, with all models successfully handling $K=2$. (c) When the turn complexity is further increased to $K=10$, performance degrades, but a clear scaling trend emerges. (d) This trend is explicitly shown, illustrating that for complex turns, the horizon length increases consistently with model size, reinforcing the benefits of scaling model size even when thinking is enabled.
@@
Appendix F Deconstructing Errors in Retrieve-then-compose Steps
@@
From Figure˜15, we can observe that models achieve near-perfect performance on the stateless retrieval and addition task, indicating that neither simple dictionary lookup nor addition is a significant source of error. In contrast, the prefix sum task, while significantly better than our task, still exhibits a slow degradation over time.
@@
Appendix G Experimental Setup
G.1 Task Details
@@ G.2 Prompting
G.3 Prompting for Thinking Models
@@ G.4 Model Specifications
We use a temperature of $0.6$ and a top-p value of $0.95$ for all Gemma models. For Qwen, we use a temperature of $0.6$ and a top-p value of $0.95$ for thinking mode and a temperature of $0.7$ and a top-p value of $0.8$ for non-thinking as recommended in their documentation. [^2] https://huggingface.co/Qwen/Qwen3-32B We find in Figure˜14 that temperature does not affect the observed trends by much.
G.5 Compute Details
@@
Appendix H Format Following Failures
@@
Our results for format following failures are presented in Figure˜16 for the experiments presented in Section˜3. We observe that smaller models are more susceptible to format failures, with the Qwen3 family in particular being worse at following format instructions. Overall, the fraction for format following errors is low (around 0.1), with the Qwen3-8B being an exception. We find that the error here actually comes from the model trying to cheat, and do the entire summation inside the <answer> tags (For example, <answer>39 + 51 = 90</answer>). This is explicitly forbidden as we do not allow chain-of-thought or thinking in this experiment, and thus we count this as an error. Gemma3 4B fails at later turns due to context length limitations; however, that does not affect any results, as its accuracies drop much earlier.
@@
For the experiments presented in Figure˜16, we find the Qwen3 family to be prone to format following errors in the case where we have thinking disabled for $K=2$. We again find this to be the consequence of models trying to cheat and use extra tokens for computation inside the answer tags. This is fixed by enabling thinking. Following this, the errors in format become negligible. At $K=10$, we see Gemma3 12B sharply rise to a format failure fraction of 1.0, again due to a full context window.
Appendix I Chain-of-Thought Self-Conditioning
@@
The alternative is to programmatically generate flawed CoT traces. We implemented and experimented with this, and as seen in Figure˜17, CoT does not mitigate the self-conditioning effect. However, this setup also introduces its own complexities. For our simple task, there are multiple distinct points of failure within a single trace: an error in the retrieval step (looking up an incorrect value) or an error in the composition step (an arithmetic mistake). A controlled experiment would need to systematically manage the type, frequency, and location of these injected errors, making the setup intractable. Even establishing a “perfectly correct” (Induced Error Rate = 0.00) baseline history is problematic. A model might have a CoT trace with flawed reasoning (e.g., a minor calculation error that cancels out), which we then replace with the correct final answer. Such a history is also unfaithful.
@@
Appendix J Proof and Analysis of Proposition 1
@@
Assuming an independent and constant per-step accuracy $p$ and no self-correction, the horizon-length $H$ at which a model can achieve a success rate $s$ is given by:
@@
J.1 Implications for Horizon Length ($H_{0.5}$)